Prove that
(1-tan^2(theta÷2))(1-tan^2(theta÷2square))(1-tan^2(theta÷2cube)).... Infinity= theta dot cot(theta)
1
Expert's answer
2019-10-31T06:26:12-0400

To show: "(1-tan^2\theta/2)(1-tan^2\theta/2^2)(1-tan^2\theta/2^3)......" "=\theta*cot\theta"


or Equivalently; "lim_{n \to \infty } (1-tan^2\theta/2)(1-tan^2\theta/2^2)..." "(1-tan^2\theta/2^n)=\theta*cot\theta"


LHS "= lim_{n \to \infty}(1-tan^2\theta/2)(1-tan^2\theta/2^2)...""(1-tan^2\theta/2^n)*tan\theta*cot\theta"

"=cot\theta *lim_{n \to \infty}\cancel{(1-tan^2\theta/2)}(1-tan^2\theta/2^2)..." "(1-tan^2\theta/2^n)(2tan\theta/2)/\cancel{(1-tan^2\theta/2)}"

"=cot\theta* lim_{n \to \infty}(1-tan^2\theta/2^2)...(1-tan^2\theta/2^n)" "(2tan\theta/2)"

"=cot\theta* lim_{n \to \infty}\cancel{(1-tan^2\theta/2^2)}...(1-tan^2\theta/2^n)" "(2^2tan\theta/2^2)/\cancel{(1-tan^2\theta/2^2)}"

"..."

"=cot\theta*lim_{n \to \infty}\cancel{(1-tan^2\theta/2^n)}" "(2^ntan\theta/2^n)/\cancel{(1-tan^2\theta/2^n)}"

"=cot\theta*lim_{n \to \infty}2^ntan\theta/2^n*(\theta/\theta)"

"=\theta*cot\theta*lim_{n \to \infty}(tan(\theta/2^n))/(\theta/2^n)"

"=\theta cot\theta"

"=RHS"


Hence Proved.


Formulas used:

  1. "tan2\theta=2tan\theta/(1-tan^2\theta)"
  2. "lim_{x \to 0} (tanx)/x =1"